OpenStudy (anonymous):

find all real solutions of each equation |2x + 1| = 6x – 9

OpenStudy (anonymous):

Split the part that equals the absolute value into plus or minus 1.) 2x+1=6x-9 and 2.) 2x+1=(-6x-9) Now just solve for x for both equations.

OpenStudy (anonymous):

for this problems I always suggest a graphic method

hero (hero):

http://www.wolframalpha.com/input/?i=solve%28 |2x+%2B+1|+%3D+6x+%E2%80%93+9%2Cx%29 Click "show steps" when you get there

hero (hero):

Make sure to copy and paste the entire url

OpenStudy (anonymous):

just solve : 2x+1=6x-9 thats the only answer

hero (hero):

yup